1) If MPC = 0.6 (and there are no income taxes) when G increases by 200, then the IS curve for any given interest rate shifts to the right by: A) 200. B) 300. C) 400. D) 500.

2)If MPC = 0.6 (and there are no income taxes but only lump-sum taxes) when T decreases by 200, then the IS curve for any given interest rate shifts to the right by: A) 100. B) 200. C) 300. D) 40
Answer #1

1) D) 500
According to multiplier formula,
Change in IS/Change in G = 1/(1-MPC) = 1/(1-0.6) = 1/0.4
So, Change in IS = Change in G*(1/0.4) = (200/0.4) = 500

2) C) 300
According to multiplier formula,
Change in IS/Change in T = -MPC/(1-MPC) = -0.6/(1-0.6) = -0.6/0.4
So, Change in IS = Change in T*(-0.6/0.4) = (-200)*(-0.6/0.4) = 300